OpenStudy (anonymous):

Will give medals! and fan! Quick help please! which of the following expressions is a sixth root of unity? A. -1 B. i C. -i D. all of the above

OpenStudy (anonymous):

@cwrw238 @hartnn

hartnn (hartnn):

what is \((-1)^6 =...\) \((-i)^6 =...\) \((i)^6 =...\) can you answer those?

OpenStudy (anonymous):

-1? yeah i feel i can

hartnn (hartnn):

whichever option gives you the answer of 1, is the sixth root of unity :)

OpenStudy (anonymous):

is that 1 and -1?

hartnn (hartnn):

in real domain, yes.

OpenStudy (anonymous):

so my only otion im given is -1, would it be accurate to choose that?

OpenStudy (anonymous):

option* sorry

hartnn (hartnn):

yes, -1 is indeed correct :)

OpenStudy (anonymous):

thank you so much!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

if i post another quick question, could you help me there?

hartnn (hartnn):

sure :)

OpenStudy (decentnabeel):

can you expal why -1 @hartnn

hartnn (hartnn):

because (-1)^6 gives you 1 (unity)

OpenStudy (decentnabeel):

but 1 is not in the option @hartnn

hartnn (hartnn):

we needed 6th root of unity let that be x so \(\sqrt[6] 1= x \\ 1 =x^6\) we know (-1)^6 = 1 so x= -1 -1 is the answer

OpenStudy (decentnabeel):

but (-1)^6=1 not -1 because the power is even so that negnative number convert in possitive.. so that why answer is -1 i donot understand @hartnn

hartnn (hartnn):

unity means 1 we both agree that (-1)^6=1 taking 6th root on both sides, -1 = sixth root of unity hence -1 is the answer.